OpenStudy (anonymous):

derterminate the number k so that the polynomial can be writen as q(x) * (x+2) and determinate q(x) x^4-4x^2-2x+k

jimthompson5910 (jim_thompson5910):

If x+2 is a factor, then x = -2 is a solution to x^4-4x^2-2x+k = 0 So simply plug in x = -2 and solve for k

OpenStudy (anonymous):

no you use long devision

jimthompson5910 (jim_thompson5910):

you can it you want, but this way is much easier

OpenStudy (anonymous):

I know its bad but I cant solve it using that T_T

jimthompson5910 (jim_thompson5910):

just replace x with -2 and simplify

jimthompson5910 (jim_thompson5910):

then solve for k

OpenStudy (anonymous):

that leads to -28 wish is not correct T_T

jimthompson5910 (jim_thompson5910):

x^4-4x^2-2x+k = 0 (-2)^4-4(-2)^2-2(-2)+k = 0 16-4(4)-2(-2)+k = 0 16-16+4+k = 0 4+k = 0 k = -4

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

np